OpenStudy (anonymous):

A ball is thrown vertically upward with an initial velocity of 80 feet per second. The distance s (in feet) of the ball from the ground after t seconds is s=80t-16t^2. a) at what time t will the ball stick the ground? b) for what time t is the ball more then 64 feet above the ground?

OpenStudy (anonymous):

when the ball hits the ground s=0

OpenStudy (anonymous):

80t-16t^2=0

OpenStudy (anonymous):

do u know how to solve that equation?

OpenStudy (anonymous):

do u put it in the quadratic equation??

OpenStudy (anonymous):

\[16t(5-t)=0\]

OpenStudy (anonymous):

u take the common factor...

OpenStudy (anonymous):

so either t=0 or t=5

OpenStudy (anonymous):

\[\therefore t=5\sec\]

OpenStudy (anonymous):

so that is the answer for part (a)

OpenStudy (anonymous):

if u got that one so i can move to solve part b

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

did u understand the first part?

OpenStudy (anonymous):

u don't sound like u did

OpenStudy (anonymous):

yeah u do 16t = 0 and (5-t)=0 then find for t

OpenStudy (anonymous):

ok good, so i did not choose t=0 to be the correct timing why?

OpenStudy (anonymous):

anyways let's do part b where u were asked to find the time at s=64ft

OpenStudy (anonymous):

64=80t-16t^2.

OpenStudy (anonymous):

do u know how to solve this one?

OpenStudy (anonymous):

16t^2-80t+64=0

OpenStudy (anonymous):

divide both sides by 16

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

not sure why you would divide 16

OpenStudy (anonymous):

to simplify the solution

OpenStudy (anonymous):

because u always want the coefficient of t^2 to be equal to 1

OpenStudy (anonymous):

okay but how will u divide 0 from 16

OpenStudy (anonymous):

16t^2-80t+64=0 ->t^2-5t+4=0

OpenStudy (anonymous):

0/16=0

OpenStudy (anonymous):

take the common factor which ever way u want to look at it all the roads lead to Rome...

OpenStudy (anonymous):

either t=1 or t=4

OpenStudy (anonymous):

okay. so the answer would be 1<x<4

OpenStudy (anonymous):

you mean t

OpenStudy (anonymous):

yes t!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if t=0 that means the maximum height is 64ft

OpenStudy (anonymous):

did u get the answer?

OpenStudy (anonymous):

so the answer would be 1<t< 4 right for b

OpenStudy (anonymous):

i replied to u but it seems my msg was not posted, yes you are right